NPP Projection Admin 3/23
To do this evening/ was completed:
Presentation Issues:
Preshow Clips/Playlist:
Trailer Updates:
Special Events/ AV set up: Tech'd the updated REC feature file for tomorrow night and turned the volume up (it looks and sounds great otherwise but was a bit too hushed at 5.5)
Presentation Issues:
Preshow Clips/Playlist:
Trailer Updates:
Special Events/ AV set up: Tech'd the updated REC feature file for tomorrow night and turned the volume up (it looks and sounds great otherwise but was a bit too hushed at 5.5)
Looked at Paper Moon for my own amusement and to see if that guest from yesterday's argument was founded in literally anything at all -- it was not, the DCP is fine, as is the pre-show content -- and also got eyes on Zola before the first round in 3
Double checked the pre-show for Conquest on 3/25 and all necessary elements are present
Ordering/Receiving (posters & trailers):
Ordering/Receiving (content): Transferred the updated REC file to theater 1 from the LMS and added it to the show playlist
Outgoing/Shipping (content): Removed Paper Moon and Zola content from theaters 1 and 3 following their shows this afternoon
Print/video Inspection:
Asset tracking doc updated/checked?:
Maintenance/cleaning: Swept and mopped in 4/6
Cue sheets updated/checked (every shift): Double checked all cues for tomorrow/cross referenced the schedule and initialed for confirmation
Marquee updated?:
Posters updated?:
Cleaning lights on for the night?:
Projection cart by entrance?:
Additional Notes: Noticed LampLOC X and Y sensor alarms in theater 1 -- no issue with playback in this room but it's something to keep an eye on as we enter April and the festival programming
Ordering/Receiving (posters & trailers):
Ordering/Receiving (content): Transferred the updated REC file to theater 1 from the LMS and added it to the show playlist
Outgoing/Shipping (content): Removed Paper Moon and Zola content from theaters 1 and 3 following their shows this afternoon
Print/video Inspection:
Asset tracking doc updated/checked?:
Maintenance/cleaning: Swept and mopped in 4/6
Cue sheets updated/checked (every shift): Double checked all cues for tomorrow/cross referenced the schedule and initialed for confirmation
Marquee updated?:
Posters updated?:
Cleaning lights on for the night?:
Projection cart by entrance?:
Additional Notes: Noticed LampLOC X and Y sensor alarms in theater 1 -- no issue with playback in this room but it's something to keep an eye on as we enter April and the festival programming
Two closed captions devices were brought down to theater 3 during the first round...open captions...screening of Black Bag and were returned to storage after the show ended
Organized letters for brunch shows (Hanna, Westward the Women) next weekend and left them behind 3/7
Comments
Post a Comment